OpenStudy (anonymous):

Solve. 8 – 3x – x = 12 x =

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

btw thx for ur help so far

OpenStudy (danjs):

8-(3+1)x = 12

OpenStudy (danjs):

8 - 4x = 12 -4x = 12 - 8 -4x = 4 x = -1